OpenStudy (anonymous):

A function is described by the equation y=2x2 - 5x-3, in which y is dependent on x. If a value for the independent variable is selected from the set {-3,0,1,3,4}, which of the following is a corresponding dependent value?

OpenStudy (er.mohd.amir):

put one by one x=-3,0,1,3,4 one i tell U y=2x2-5x-3 y=2(-3)2-5(-3)-3 =2*9+15-3=18+15-3=30

OpenStudy (er.mohd.amir):

y=2(0)^2-5(0)-3=0-0-3=-3